OpenStudy (anonymous):

the profit that the vender makes per day by selling x pretzels is given by the function p(x)=-0.004x^2+2.8x-250. find the number of pretzels that must be sold to maximize profit

OpenStudy (austinl):

Do you have a graphing calculator?

OpenStudy (austinl):

Are you familiar with the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Wait... nevermind, that wont work.

OpenStudy (austinl):

We have a function of the form, \(p(x)=ax^2+bx+c\) We can find the maximum y(\(p(x)\)) value using this formula. \(\displaystyle p(x)_{max/min}=c-\frac{b^2}{4a}\) We can plug in values for that, and then when we figure that out, we can solve for x at that point!